A particle moves along a straight line with constant acceleration a=3 m/s2


. At t=0 the


position of the particle is x0=3 m and the particle is at rest (v0=0).


i) Find the position and the velocity of the particle at t=2 s [3 mks]


ii) Find the position of the particle when v=30 m/s [3 mks]

1
Expert's answer
2021-12-14T09:27:19-0500

The information given is

The initial position is "X_{\circ}=0\;m"

The initial velocity is "V_{\circ}=0 \; m/s"

The acceleration is "a=3\;m/s^{2}"


Part i

The position is given by

"X=X_{\circ}+V_{\circ}\;t+\dfrac{1}{2}\;a\;t^{2}"


Evaluating numerically in t=2 s


"X=X_{\circ}+V_{\circ}\;t+\dfrac{1}{2}\;a\;t^{2}\\
X=0\;m+0\;m/s\times 2\;s+\dfrac{1}{2}\times 3\;m/s^{2}\times (2\;s)^{2}\\
X=6\;m"


The position at t-2 s is "X=6\;m"


The velocity is given by

"V=V_{\circ}+a\;t"

Evaluating numerically.

"V=V_{\circ}+a\;t\\
V=0\;m/s+3\;m/s\times 2\;s\\
V=6\;m/s"


The velocity at t= 2s is "V=6\;m/s"


Part ii

The velocity is given by

"V=V_{\circ}+a\;t"

Obtaining the expression for the time.

"V=V_{\circ}+a\;t\\\\
V_{\circ}+a\;t=V\\\\
a\;t=V-V_{\circ}\\\\
t=\dfrac{V-V_{\circ}}{ a}\\\\"

Evaluating numerically.

"t=\dfrac{30\;m/s-0\;m/s}{ 3\;m/s^{2}}\\\\
t=10\;s"

The time where the velocity is 30 m/s is "t=10\;s"


The position is given by 

"X=X_{\circ}+V_{\circ}\;t+\dfrac{1}{2}\;a\;t^{2}"

Evaluating numerically.

"X=0\;m+0\;m/s\times 10\;s+\dfrac{1}{2}\times 3\;m/s^{2}\times (10\;s)^{2}\\
X=150\;m"


The position when V=30 m/s is "X=150\;m"